A plane travels at a speed of 205 mph in still air. Flying with a tailwind, the plane is clocked over a distance of 800 miles. Flying against a headwind, it takes 1 hour more time to complete the return trip. What was the wind velocity?
Answer by mananth(16946)

plane speed = 205 mph
wind sped = x mph
against wind = 205-x
with wind = 205 +x
t=d/r
time against wind - time with wind =1
800/(205-x)-800/(205+x) =1
800(205+x)-800(205-x)=(205-x)(205+x)
164000+800x-164000+800x=42025-x^2
1600x+x^2-42025=0
solve for x the wind speed should get wind speed = 25.85 mph